### v2

# First Light and Reionization Epoch Simulations (FLARES) - XV: The physical properties of super-massive black holes and their impact on galaxies in the early Universe

Abstract
Understanding the co-evolution of super-massive black holes (SMBHs) and their host galaxies remains a key challenge of extragalactic astrophysics, particularly the earliest stages at high-redshift. However, studying SMBHs at high-redshift with cosmological simulations, is challenging due to the large volumes and high-resolution required. Through its innovative simulation strategy, the First Light And Reionisation Epoch Simulations (FLARES) suite of cosmological hydrodynamical zoom simulations allows us to simulate a much wider range of environments which contain SMBHs with masses extending to $M_{\bullet}>10^{9}\ {\rm M_{\odot}}$ at $z=5$. In this paper, we use FLARES to study the physical properties of SMBHs and their hosts in the early Universe ($5\leq\,z\leq 10$). FLARES predicts a sharply declining density with increasing redshift, decreasing by a factor of 100 over the range $z=5\to 10$. Comparison between our predicted bolometric luminosity function and pre-JWST observations yield a good match. However, recent JWST observations appear to suggest a larger contribution of SMBHs than previously observed, or predicted by FLARES. Finally, by using a re-simulation with AGN feedback disabled, we explore the impact of AGN feedback on their host galaxies. This reveals that AGN feedback results in a reduction of star formation activity, even at $z>5$, but only in the most massive galaxies. A deeper analysis reveals that AGN are also the cause of suppressed star formation in passive galaxies but that the presence of an AGN doesn’t necessarily result in the suppression of star formation.
1 Introduction
Since the first conceptualization of black holes almost 250 years ago by John Mitchell and Pierre–Simon Laplace (see e.g. Schaffer, 1979; Montgomery et al., 2009) and the first observation of a quasar 60 years ago by Schmidt (1963), numerous scientists have worked on understanding how black holes form, evolve, and affect their surroundings (some of the most seminal works include Kerr, 1963; Salpeter, 1964; Penrose, 1965; Lynden-Bell, 1969; Penrose & Floyd, 1971; Bardeen et al., 1973; Shakura & Sunyaev, 1973; Blandford & Znajek, 1977; Abramowicz et al., 1988; Narayan & Yi, 1994). However, a complete theory of how black holes operate and interact with their host galaxies (e.g. Rees, 1984; Richstone et al., 1998) remains one of the biggest challenges in modern (astro)physics today.
[…]
Observations of high redshift quasars (e.g. Jiang et al., 2016; Matsuoka et al., 2016; Maiolino et al., 2023b) have revealed that SMBHs existed in the Universe less than a billion years after the Big Bang (see Inayoshi et al., 2020; Fan et al., 2023, for recent reviews). The traditional stellar remnant BH formation scheme cannot explain such massive BHs in the early Universe; in order to grow BHs up to $\sim 10^{9}\ {\rm M_{\odot}}$ at $z\gtrsim 5$, alternative formation mechanisms are required, such as massive seeds and/or enhanced BH accretion (Latif et al., 2013; Volonteri et al., 2021). Suggested massive seed formation scenarios include direct collapse black holes, remnants of Population III stars, and the collapse of very massive stars formed through mergers (Loeb & Rasio, 1994; Madau & Rees, 2001; Bromm & Loeb, 2003; Portegies Zwart et al., 2004; Volonteri & Rees, 2005; Begelman et al., 2006; Regan & Haehnelt, 2009). Understanding not only the formation but also the evolution and physics of high redshift SMBHs is essential in order to capture the effects that SMBHs have on their surroundings. Since SMBHs grow by accreting surrounding gas while simultaneously releasing energy into it (e.g. Fabian, 2012), they affect their surroundings, thus altering the overall properties of their host galaxies (e.g. the bright end of the galaxy luminosity function Kauffmann & Haehnelt, 2000; Granato et al., 2004; Bower et al., 2006; Cattaneo et al., 2006; Croton et al., 2006).
In addition to influencing its host galaxy, the radiation emitted in the vicinity of black holes also contributes to the overall ionizing photon budget, although stellar sources of ultraviolet photons seem to predominantly drive the hydrogen reionization of the Universe (Madau & Haardt, 2015; Qin et al., 2017; Dayal & Ferrara, 2018; Robertson, 2022). Since the number density of Active Galactic Nuclei (AGN) increases rapidly towards lower redshifts, the fractional contribution of AGN to the total ionizing photon budget becomes more significant towards lower redshifts. Although their ionizing photon contribution during hydrogen reionization was not dominant, their higher number density and harder spectra suggest that AGN significantly contributed to helium reionisation and to the meta-galactic UV and X-ray background of the Universe (e.g. Ricotti & Ostriker, 2004; Giallongo et al., 2019; Puchwein et al., 2019; Finkelstein & Bagley, 2022).
From a theoretical/computational point of view, black hole physics has been an integral component of models of galaxy formation, which try to capture the effects of black hole feedback on the simulated galaxies (see the reviews of Somerville & Davé, 2015; Naab & Ostriker, 2017; Vogelsberger et al., 2020; Habouzit et al., 2022a, b). Since SMBHs grow by accreting surrounding gas while simultaneously releasing energy to it (e.g. Fabian, 2012), they affect their surroundings thus altering the overall properties of their host galaxies. Traditionally, black hole feedback has been incorporated either through a thermal / quasar mode, where a fraction of the bolometric luminosity is injected as thermal energy to the surrounding environment (Springel et al., 2005; Booth & Schaye, 2009; Tremmel et al., 2017) or as kinetic (also known as mechanical or radio mode) mode (Croton et al., 2006; Costa et al., 2014; Choi et al., 2015; Costa et al., 2020), or as a combination of different modes (Sijacki et al., 2007; Dubois et al., 2012; Sijacki et al., 2015; Weinberger et al., 2017; Davé et al., 2019). However, different implementations of black hole physics result in discrepancies in the predictions of black hole properties both at low and at high redshifts (Meece et al., 2017; Habouzit et al., 2022a, b), which makes understanding the co–evolution of black holes and galaxies even more challenging.
With the advent of JWST the observational SMBH frontier is now shifting to higher-redshift. Samples of SMBHs have now been detected out to $z\approx 10$, deep into the Epoch of Reionisation (Larson et al., 2023; Harikane et al., 2023; Juodžbalis et al., 2023; Matthee et al., 2023; Greene et al., 2023; Kocevski et al., 2023; Maiolino et al., 2023b; Übler et al., 2023; Kokorev et al., 2024) with tentative detections at $z>10$ (Maiolino et al., 2023a; Bogdán et al., 2023; Juodžbalis et al., 2023). The innovation of JWST is its ability to constrain AGN activity in galaxies through broad line emission, line-ratios, compact morphology, broad-band photometry, or a combination thereof. With new imaging and spectroscopic surveys underway, or planned, samples of high-redshift AGN will inevitably grow in size and robustness. JWST observations will also soon be complemented by wide-area observations from Euclid, providing large samples of bright, AGN-dominated sources.
The contribution of JWST, and soon Euclid, represents an important new frontier in cosmological galaxy formation. Comparison between these observations and galaxy formation models will provide the opportunity to constrain the formation and growth mechanisms of SMBHs in the early Universe.
However, simulating large samples of SMBH dominated galaxies in the early Universe is challenging due to their relative rarity, thus requiring large simulations. Flagship simulations such as Illustris (Vogelsberger et al., 2014b, a; Genel et al., 2014; Sijacki et al., 2015), EAGLE (Schaye et al., 2015; Crain et al., 2015; McAlpine et al., 2016, 2017), Horizon-AGN (Dubois et al., 2016; Volonteri et al., 2016), TNG100 (Weinberger et al., 2017; Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018b, a; Springel et al., 2018; Weinberger et al., 2018), Simba (Davé et al., 2019), etc., are too small to yield statistically useful samples of observationally accessible massive SMBHs in the early Universe (see Habouzit et al., 2022b). While larger simulations exist, including BAHAMAS (McCarthy et al., 2017), TNG300 (Weinberger et al., 2017; Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2018; Pillepich et al., 2018b, a; Springel et al., 2018; Weinberger et al., 2018), FLAMINGO (Schaye et al., 2023), most have significantly lower mass-resolution, limiting their use for studying the SMBHs now accessible to JWST. The exceptions are simulations that only target the high-redshift Universe, for example: Massive Black (Khandai et al., 2012), Bluetides (Di Matteo et al., 2017; Wilkins et al., 2017; Tenneti et al., 2018; Huang et al., 2018; Ni et al., 2020; Marshall et al., 2020), ASTRID (Bird et al., 2022; Ni et al., 2022), and more recently the First Light And Re-ionisation Epoch Simulations (FLARES, Lovell et al., 2021; Vijayan et al., 2021), the focus of this study.
[…]
This paper is structured as follows: in Section 2 we detail the simulation suite FLARES as well as modelling methodology for SMBH and stellar emission. In Section 3 we present predictions for the physical and observational properties of SMBHs, including the environmental dependence (§3.2.3). In Section 4 we explore the correlation of SMBH properties with the properties of their hosts. In Section 4.4 we briefly discuss the impact of SMBHs on their host galaxies, and finally in Section 5 we present our conclusions.
2 Simulations and Modelling
[…]
2.1 FLARES
Figure 1: The number of $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ (thin line) and $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ (thick line) super–massive black holes in FLARES (solid–line) and the EAGLE (dashed line) reference volume at $z=5-10$.
FLARES is a suite of 40 hydrodynamical re-simulations of spherical regions of size $\rm 14\,cMpc\,$$h^{-1}$ utilising the AGNdT9 variant of the EAGLE (Schaye et al., 2015; Crain et al., 2015) physics model and a Planck year 1 cosmology ($\rm\Omega_{M}=0.307,\;\Omega_{\Lambda}=0.693,\;$h$\;=0.6777$, Planck Collaboration et al., 2014). The regions were selected from a large (3.2 cGpc)${}^{3}$ low-resolution dark matter only (DMO) simulation (Barnes et al., 2017). The selected regions encompass a wide range of overdensities, $\delta_{14}=-0.4\to 1.0$, with greater representation at the extremes, particularly extreme over-densities. This enables us to simulate many more massive galaxies than possible using a periodic simulation and the same computational resources. For example, FLARES contains approximately 100 times as many $M_{\star}>10^{9}\ {\rm M_{\odot}}$ galaxies at $z=10$ than the fiducial EAGLE reference simulation. This makes FLARES ideally suited to studying SMBHs, and particularly AGN, since they are rare and preferentially occur in massive galaxies. This is demonstrated in Figure 1, where we show the total number of SMBHs with $M_{\bullet}/{\rm M_{\odot}}>\{10^{7},10^{8}\}$ as a function of redshift for both the (100 Mpc)${}^{3}$ EAGLE reference simulation and all FLARES regions combined. At $z=5$ FLARES contains 6 (20) times as many SMBHs with $M_{\bullet}>10^{7}(10^{8})\ {\rm M_{\odot}}$, respectively. This disparity increases with increasing redshift: there are no SMBHs with $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ in EAGLE at $z>6$, for example, whereas those in FLARES extend to $z=9$.
2.1.1 Weighting
An important consequence of the FLARES strategy is that universal cosmological scaling relations and distribution functions (e.g. the SMBH mass function) cannot be trivially recovered. Instead, it is necessary to weight each simulation/region by how likely it is to occur in the parent DMO simulation. This is described in more detail in Lovell et al. (2021) where we demonstrate its appropriateness by recovering the galaxy stellar mass function. In brief, we measure the overdensity in the parent box on a $2.67\;\mathrm{Mpc}$ grid, find the distribution of overdensities within each region, and then weight each region $j$ by some factor
$f_{j}=\sum_{i}r_{i}w_{ij}\;\;,$ (1)
where $r_{i}$ is the ratio of the true weighting in overdensity bin $i$ (over the whole parent box) to the summed weight of all regions in that overdensity bin, and $w_{ij}$ is the weighting of region $j$ in overdensity bin $i$.
2.2 Black Hole Modelling in EAGLE
In this section we summarise the SMBH physics utilised by the EAGLE model which is employed by FLARES. For a full description of FLARES see Lovell et al. (2021); Vijayan et al. (2021), and for the EAGLE model see Schaye et al. (2015) and Crain et al. (2015). In short, BHs in the EAGLE model are seeded into sufficiently massive halos and then allowed to grow through accretion and mergers. A fraction of the rest-mass accreted onto the disc is able to be radiated away. A fraction of radiated energy is injected into neighbouring gas particles, heating them.
2.2.1 Seeding
BHs are seeded into halos exceeding a halo mass of $\log_{10}(M_{\rm h}/h^{-1}{\rm M_{\odot}})=10$ by converting the highest density gas particle into a BH particle (Springel et al., 2005). BHs carry both a particle mass and a subgrid mass. The particle initial mass is set by that of the converted gas particle while the initial subgrid mass is $\log_{10}(M_{\bullet}/h^{-1}{\rm M_{\odot}})=5$. The use of a separate subgrid mass is necessary because the black hole seed mass is below the simulation mass resolution. Calculations pertaining to growth and feedback events of the black hole are computed using the subgrid mass, $M_{\bullet}$, while gravitational interactions use the particle mass. When the sub-grid BH mass exceeds the particle mass the BH particle is allowed to stochastically accrete a neighbouring gas particle. When the sub-grid BH mass is much larger than the gas particle mass, the BH sub-grid and particle masses effectively grow together. The seed and gas particle masses place an effective lower-limit on the BH masses which are robust, or resolved. For this analysis we conservatively assume that BHs are considered resolved at $M_{\bullet}=10^{7}\ {\rm M_{\odot}}$ and focus our analysis on these objects.
2.2.2 Accretion
The primary growth of BHs is through accretion. The EAGLE subgrid accretion model allows the BH to accrete material at a maximum rate determined by the Eddington accretion rate scaled by a factor of $1/h$ (McAlpine et al., 2020), i.e.
$\dot{m}_{\rm Edd}=\frac{4\uppi Gm_{\bullet}m_{\rm p}}{\epsilon_{\rm r}\sigma_{\rm T}c},$ (2)
where $G$ is the gravitational constant, $m_{\bullet}$ is the black hole subgrid mass, $m_{\rm p}$ is the proton mass, $\epsilon_{\rm r}$ is the radiative efficiency of the accretion disk, $\sigma_{\rm T}$ is the Thomson cross-section and $c$ is the speed of light. The radiative efficiency is set to $\epsilon_{\rm r}=0.1$.
The floor for the accretion rate is set by the Bondi–Hoyle accretion rate defined as
$\dot{m}_{\rm Bondi}=\frac{4\uppi G^{2}m_{\bullet}^{2}\rho}{(c_{\rm s}^{2}+v^{2})^{3/2}},$ (3)
where $c_{\rm s}$ and $v$ are the speed of sound and the relative velocity of the BH and gas, respectively (Bondi & Hoyle, 1944). Hence, the actual accretion rate onto the black hole is then given by
$\dot{m}_{\rm accr}=\dot{m}_{\rm Bondi}\times\mathrm{min}\left(C_{\rm visc}^{-1}(c_{\rm s}/V_{\upphi})^{3},1\right),$ (4)
where $C_{\rm visc}$ is a parameter related to the viscosity of the accretion disc (see next subsection) and $V_{\upphi}$ is the rotation speed of the gas around the BH (see equation 16 of Rosas-Guevara et al., 2015). To re-iterate, the maximum of the accretion rate is set by equation 2. Finally, this results in the BH mass growth rate of
$\dot{m}_{\rm\bullet}=(1-\epsilon_{\rm r})\dot{m}_{\rm accr}.$ (5)
The accretion rates of each BH are reported at every time–step the BH is active and, for all BHs, at the snapshot redshift. The rates reported in a single time–step can differ significantly from rates averaged across longer timescales. In Appendix A we contrast the instantaneous accretion rates with those averaged on longer timescales.
2.2.3 Emission
The total (bolometric) luminosity radiated by an AGN is simply proportional to the accretion disc or BH growth rate as,
$L_{\rm\bullet,bol}=\epsilon_{\rm r}\dot{m}_{\rm accr}c^{2}=\left(\frac{\epsilon_{\rm r}}{1-\epsilon_{\rm r}}\right)\dot{m}_{\bullet}c^{2}.$ (6)
As noted previously, the radiative efficiency $\epsilon_{\rm r}$ is assumed to be $0.1$ in the EAGLE model.
In this work we only present this bolometric luminosity. However, in a companion work (Wilkins et al. in-prep) we fully model the spectral energy distributions, including the contribution of nebular continuum and line emission, of AGN using a subgrid model combining the disc, narrow and broad line regions, and a dusty torus.
2.2.4 Mergers
In addition to accretion, BHs can also grow by merging. BHs are merged if their separation is smaller than three gravitational softening lengths and are within the BH smoothing kernel length, $h_{\bullet}$, of each other and have a relative velocity of
$v_{\mathrm{rel}}<\sqrt{Gm_{\bullet}/h_{\bullet}},$ (7)
where $m_{\bullet}$ is the subgrid mass of the bigger of the merging BHs. The limiting velocity given by equation 7 prevents BH mergers during the initial stages of galaxy mergers and makes them only possible once some time of the initial merger has passed and the relative velocities have settled. As a consequence galaxies in the EAGLE model can often host multiple black holes. We explore this in the context of FLARES in §3.1.
2.2.5 Repositioning
[…]
2.2.6 Feedback
In terms of the subgrid AGN feedback model FLARES adopts the modelling approach of the C-EAGLE simulations (Barnes et al., 2017), using the AGNdT9 subgrid parameter configuration. This configuration is parameterised with $C_{\mathrm{visc}}=2\uppi\times 10^{2}$ and $\Delta T_{\rm AGN}=10^{9}\,$, where the former is a free parameter controlling the sensitivity of the BH accretion rate to the angular momentum of the gas and the latter is the temperature increase of gas during AGN feedback.
AGN feedback in the EAGLE model is modelled with only one feedback channel, contrasting with the multi-mode feedback implemented in other simulations including Simba (Davé et al., 2019) and TNG (Zinger et al., 2020). In this model, thermal energy is injected stochastically to gas particles neighbouring the BH particle, in a kernel–weighted manner. The energy injection rate is calculated as a fraction of the total accretion rate, as $\epsilon_{\rm f}\epsilon_{\rm r}\dot{m}_{\rm accr}c^{2}$, where $\epsilon_{\rm f}=0.15$ is the fraction of the feedback energy that is coupled to the ISM and $\epsilon_{\rm r}=0.1$ is the radiative efficiency of the accretion processes introduced in Eq. 5.
The probability of injecting energy into each nearby gas particle is given by
$P=\frac{E_{\bullet}}{\Delta\epsilon_{\rm AGN}N_{\mathrm{ngb}}\left\langle m_{\rm g}\right\rangle},$ (8)
where $E_{\bullet}$ represents a ‘reservoir’ of feedback energy carried by the black hole, with $\epsilon_{\rm r}\epsilon_{\rm f}\dot{m}_{\rm accr}c^{2}\Delta t$ added to it after every time step $\Delta t$, $\Delta\epsilon_{\rm AGN}$ is the change in internal energy per unit mass of a gas particle corresponding to a temperature increase of $\Delta T_{\rm AGN}$, $N_{\mathrm{ngb}}$ is the number of gas neighbours of the BH and $\left\langle m_{\rm g}\right\rangle$ is their mean mass. The reservoir, $E_{\bullet}$, is then decreased by the amount of the injected energy.
Since $\Delta T_{\rm AGN}$ directly determines the amount of energy in each AGN feedback event, it is the most important factor in modelling the feedback from BH accretion. A larger value results in more energy dumped into the neighbouring particles of the BH, but also makes the feedback events more rare, as the change in internal energy of a gas particle is directly proportional to temperature increase and the probability of energy injection into a gas particle is inversely proportional to the internal energy increase. The temperature increase from AGN feedback was set to a value higher than that from stellar feedback due to gas densities being higher in the vicinity of black holes than they are for typical star-forming gas (Crain et al., 2015).
3 Blackhole Properties
In this section we explore the physical properties of SMBHs and their host galaxies in FLARES. As previously noted (in §2.2), a consequence of the simulation resolution and modelling choices (i.e. SMBH seed mass) is that we can only be confident in the properties of SMBHs with masses $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$, which we consider ”resolved”. Consequently, we focus our attention on these systems though we do explore predictions at lower-masses.
3.1 Multiplicity
As noted in §2.2.4, galaxies in the EAGLE model can host multiple black holes, though many of these will eventually merge. We find that while most massive galaxies simulated by FLARES contain multiple black holes, in the vast majority ($\approx 94\%$) of galaxies hosting at least one resolved SMBH, the most-massive SMBH accounts for $>90\%$ of the total mass. Out of the galaxies hosting a resolved SMBH, around 2% of them host multiple resolved black holes.
3.2 SMBH density and mass Function
Figure 2: The evolution of the SMBH number density predicted by EAGLE and FLARES at $z=10-5$ for $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ and $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$.
Next, we explore the evolution of the SMBH number density, from $z=10\to 5$, in Figure 2. As expected we find good agreement between FLARES and EAGLE, validating our simulation strategy and weighting scheme.
The density drops by 10$\approx$ 10$\times$ 10 × between $z=5\to 7$, irrespective of the mass threshold. This is comparable to the drop in density of galaxies with $M_{\star}>10^{10}\ {\rm M_{\odot}}$ (see Lovell et al., 2021), but faster than the drop in lower-mass galaxies. This evolution appears to continue at $z>7$ but here the numbers simulated are small, leading to a large statistical uncertainty.
Figure 3 shows the SMBH mass function predicted by FLARES, again from $z=10\to 5$. The mass function also drops by 10$\approx$ 10$\times$ 10 × from $M_{\bullet}=10^{7}\to 10^{8}\ {\rm M_{\odot}}$, largely independent of redshift. There is tentative evidence of a steeper drop at $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$, but this is complicated by the small sample size and thus large statistical uncertainty. Figure 3 also shows the SMBH mass function split by bolometric luminosity. Of note here is that the most massive SMBHs are not necessarily the most luminous; this is discussed in more detail in §3.3.2. The total mass function shown in Figure 3 is provided in Table 1 and electronically.
Figure 3: The evolution of the SMBH mass function predicted by FLARES at $5\leq z\leq 10$. The solid coloured lines in both panels show the total mass function (i.e. for all SMBHs). The shaded region around this provides an estimate of the uncertainty using the $68\%$ Poisson confidence interval. The thin coloured dashed, dot-dashed, and dotted lines on the left-hand panel denote the mass functions of SMBHs with $L_{\rm bol}/{\rm erg\ s^{-1}}<10^{45}$, $10^{45}<L_{\rm bol}/{\rm erg\ s^{-1}}<10^{46}$, and $L_{\rm bol}/{\rm erg\ s^{-1}}>10^{46}$ respectively. Shown on the right-hand panels are recent observational constraints and model predictions. The observations Matthee et al. (2023) at $z\sim 5$ and He et al. (2023) at $z\approx 4$ in the $z=5$ panel. Models predictions include ASTRID, Bluetides, EAGLE, Illustris, Simba, TNG100, and TNG300.
Table 1: The SMBH mass function at $5\leq z\leq 10$ predicted by FLARES. An electronic version of this is available at: https://github.com/stephenmwilkins/flares_agn_data.
$\log_{10}(\phi/{\rm Mpc^{-3}\ dex^{-1}})$
$\log_{10}(M_{\bullet}/{\rm M_{\odot}})$ $z=10$ $z=9$ $z=8$ $z=7$ $z=6$ $z=5$
7.25 -7.0 -5.47 -4.99 -4.87 -4.79 -4.05
7.5 -6.3 -7.54 -6.42 -4.89 -4.57 -4.15
7.75 - -6.22 -6.15 -5.98 -4.72 -4.31
8.0 - -7.32 -5.52 -6.06 -5.11 -4.73
8.25 - - -6.76 -5.48 -5.85 -4.99
8.5 - - -7.32 -6.65 -5.35 -5.24
8.75 - - - -7.32 -6.51 -5.16
9.0 - - - - -6.89 -7.32
9.25 - - - - - -6.89
3.2.1 Observational Constraints
As a key physical distribution function the SMBH mass function has been the focus of considerable observational study.
[…]
He et al. (2023) recently combined SDSS observations with a fainter Hyper Suprime-Cam selected sample with spectroscopic follow-up to study the SMBH mass function at $z\approx 4$. This samples ranges from $M_{\bullet}=10^{7.5-10.5}\ {\rm M_{\odot}}$ and bolometric luminosities $10^{45.5-47.5}\ {\rm erg\ s^{-1}}$. The broad line SMBH mass function at $z=4$ observed by He et al. (2023) is shown alongside our predictions in Figure 3. These constraints are consistent with our total mass function function at $\approx 10^{9}\ {\rm M_{\odot}}$ but fall short at lower-masses. However, the He et al. (2023) mass function is measured at $z=4$, not at $z=5$ where the FLARES predictions lie. Extrapolating the FLARES $z=5$ SMBH mass function to $z=4$ would suggest a density increase of around 0.5 dex. Secondly, the He et al. (2023) mass function only includes unobscured (i.e. type 1) and active SMBH and thus provides only a lower-limit on the true mass function. However, He et al. (2023) also attempt to constrain the total mass function, making corrections for obscuration and in-active SMBHs. The density of $M_{\bullet}=10^{8-9}\ {\rm M_{\odot}}$ BHs is 10100$\sim$ 10-100$\times$ 10 - 100 × larger than density of active broad line SMBHs, with the range sensitive to the modelling assumptions. The upper-end of this correction would elevate the total mass function above our predicted mass function across the full mass-range. While highly uncertain this suggests that the FLARES predictions are compatible with these observations.
JWST is now enabling a similar approach at higher-redshift and lower luminosities. Several studies (Harikane et al., 2023; Matthee et al., 2023; Maiolino et al., 2023b) have already employed this method to infer SMBH masses, and in the case of Matthee et al. (2023) the broad-line SMBH mass function. The observational constraints of Matthee et al. (2023) are shown in Figure 3 and, at first glance, provide a good match to our predictions for all SMBH. However, for the same reasons described above, the Matthee et al. (2023) mass function will provide a lower-limit on the total mass function. This suggests possible tension between our predictions and Matthee et al. (2023) and possibly Matthee et al. (2023) and He et al. (2023).
This comparison highlights the issues involved in directly comparing predictions from simulations with observations. These issues can be lessened, though not eliminated, by instead comparing bolometric luminosity functions (which we do in §3.3.2). However, this results in discarding useful information (the line-widths). An alternative approach is to forward model the SMBHs to predict observational quantities that can be analysed in the same way as observational samples. This is a current focus of work with results expected in a companion work (Wilkins et al. in-prep).
3.2.2 Model comparisons
The right-hand panel of Figure 3 also shows predictions from several other cosmological hydrodynamical simulations including Astrid, Bluetides, EAGLE, Illustris, Simba, TNG100, and TNG300. Figure 3 immediately demonstrates the power of the FLARES approach. Only the $(400/h\ {\rm Mpc})^{3}$ Bluetides simulation mass function extends as far as FLARES, and only then to $z=7$ where the simulation stopped. For all the other simulations FLARES significantly extends the mass range probed. As expected, the FLARES predictions closely match the EAGLE predictions (where they overlap) but are also similar to Illustris, Simba, TNG100, and Bluetides (at $z=7$, less so at $z>7$). The agreement with ASTRID and TNG300 is weaker with both predicting a lower normalisation than found in FLARES. A thorough comparison between many of these models is presented in Habouzit et al. (2022a) at low-redshift and Habouzit et al. (2022b) at high-redshift.
3.2.3 Environmental Dependence
One of the strengths of FLARES is its ability to probe the effect of environment on galaxy formation at high-redshift. In Figure 4 we show how the total number and mass function of SMBHs varies between simulations and thus environment, $\log_{10}(1+\delta_{14})\approx-0.3\to 0.3$.
These figures reveal, unsurprisingly, that SMBHs in FLARES are extremely biased. Virtually all of the $>10^{7}\ {\rm M_{\odot}}$ SMBHs simulated in FLARES are in the most extreme regions. While there are $\approx 700$ $>10^{7}\ {\rm M_{\odot}}$ SMBHs across the FLARES simulations at $z=5$, only $\approx 10$ are in regions with $\delta_{14}\leq 0.0$. The strong environmental dependence of SMBHs predicted by FLARES raises the issue of cosmic variance affecting observational surveys (e.g. Thomas et al., 2023).
Figure 4: Environmental dependence of the number (top panel) and mass function (bottom panel) of SMBHs predicted by FLARES at $z=5$. The top panel shows the number of robust $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs in each simulated region as a function of the region over-density $\delta_{14}$. Dotted lines indicate the behaviour if $N\propto\delta+1$. The bottom panel shows the SMBH mass function for individual simulations, coloured coded by the region over-density $\delta_{14}$. Also shown here in black is the composite SMBH mass function found by combining all forty simulations with the appropriate weighting. The horizontal grey line denotes the density corresponding to a single object in an individual FLARES simulation.
3.3 Growth
As described in §2.2 SMBHs in the EAGLE model grow through both accretion (§2.2.2) and mergers (§2.2.4). We defer an exploration of SMBH mergers in FLARES to a work in preparation (Liao et al. in-prep) and focus here on growth through accretion, including making predictions for the (accretion disc) accretion rates and bolometric luminosities of SMBHs in FLARES. As explained in §2.2.3, since we assume a fixed radiative efficiency (in our case $\epsilon_{\rm r}=0.1$) the bolometric luminosity simply scales with the accretion disc accretion rate, allowing us to explore them interchangeably.
3.3.1 Correlation with SMBH mass
We begin, in Figure 5, by showing predictions for the relationship between SMBH mass ($M_{\bullet}$) and the accretion rate, bolometric luminosity (top–panel) and Eddington ratio $\lambda$ (bottom–panel). As noted previously we can only be confident in our predictions of SMBHs with $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. Since a large fraction of SMBHs are accreting at the Eddington limit this suggests our predictions for the bolometric luminosities of SMBHs are, conservatively, complete above the Eddington luminosity of a $10^{7}\ {\rm M_{\odot}}$ SMBH, i.e. $\approx 10^{45}\ {\rm erg/s}$ ($\approx 2.5\times 10^{11}\ {\rm L_{\odot}}$).
SMBHs in FLARES exhibit a wide range of accretion rates at fixed mass, extending up to our imposed maximum ($1/h$$\times$ the Eddington rate). For resolved SMBHs the median Eddington ratio ($\lambda$) is around $0.01$ below which it gradually drops such that there are very few galaxies with $\lambda<10^{-6}$. While the number of SMBHs with ratios higher than the median also drops, they “pile-up” at the Eddington luminosity due to the imposed limiter. The binned median Eddington ratio (denoted by the dashed line in the bottom–panel of Figure 5) drops by around $0.5$ dex $M_{\bullet}>10^{7}-10^{9}\ {\rm M_{\odot}}$. Figure 5 also shows the binned median Eddington rate but weighted by the accretion rate (solid). As would be expected this is biased towards higher accretion rates, except at the highest masses where the small numbers lead to convergence. Thus, while the typical Eddington ratios are $\sim 0.01$ most accretion (and thus most energy is produced) is SMBHs with much higher Eddington ratios.
Another feature to note is that the most luminous SMBHs in FLARES are not necessarily the most massive. While more massive SMBHs on average have higher luminosities, the mass function is so steep that there are many more lower-mass SMBHs resulting in them making up a larger share of the most luminous SMBHs. For example, at $z=5$ the most luminous ($L_{\rm bol}\approx 10^{47}\ {\rm erg\ s^{-1}}$) SMBH has a mass of $M_{\bullet}\approx 2\times 10^{8}\ {\rm{\rm M_{\odot}}}$ compared to the most massive SMBH which has $M_{\bullet}>10^{9}\ {\rm{\rm M_{\odot}}}$.
Recent observational constraints from JWST suggest Eddington ratios of $\lambda=0.01-1.0$ in $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ (e.g. Maiolino et al., 2023b). Our predictions extend to lower values of $\lambda$ but this may simply reflect an observational bias, since SMBHs with lower ratios will be less luminous and thus possibly missed. Thus, at present, there does not appear to be a contradiction between our predictions and the observations of these properties.
Figure 5: Top: The relationship between SMBH mass and accretion disc accretion rate (left-axis) and bolometric luminosity (right-axis) predicted by FLARES at $z=5$. Diagonal lines denote fixed Eddington ratios $\lambda\in\{0.001,0.01,0.1,1.0\}$ and objects are colour-coded by their Eddington ratio. Bottom: The relationship between SMBH mass and Eddington ratio $\lambda$. The thick dashed line denotes the binned median while the solid line denotes the binned median weighted by the accretion rate. The right-hand panel shows the normalised distribution of Eddington ratios for $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. The dotted, solid, and dashed lines denote the 15.8th, 50th, and 84.2th percentiles respectively.
3.3.2 Bolometric Luminosity Function
Figure 6: Bolometric luminosity function for SMBHs at $z\in[5,10]$ predicted by FLARES. Also shown are high-redshift predictions for the SMBH bolometric luminosity function from the Shen et al. (2020) empirical models of and recent observational constraints on the SMBH bolometric luminosity function from Kokorev et al. (2024). The faint grey line in each panel shows the $z=5$ prediction.
Table 2: The SMBH bolometric luminosity function at $5\leq z\leq 10$ predicted by FLARES. An electronic version of this is available at: https://github.com/stephenmwilkins/flares_agn_data.
$\log_{10}(\phi/{\rm Mpc^{-3}\ dex^{-1}})$
$\log_{10}(L_{\rm bol}/{\rm erg\ s^{-1}})$ $z=10$ $z=9$ $z=8$ $z=7$ $z=6$ $z=5$
45.1 -5.4 -5.75 -6.63 -5.75 -5.16 -4.6
45.3 -6.75 -6.68 -6.17 -5.63 -4.76 -4.34
45.5 -6.9 -6.09 -6.38 -5.63 -5.83 -5.04
45.7 - - -6.75 -6.05 -5.81 -5.36
45.9 - -7.18 -7.28 -6.71 -7.04 -6.51
46.1 - - -7.45 -6.2 -6.34 -6.36
46.3 - -7.22 -5.42 -7.45 -6.29 -4.98
Next, in Figure 6, we explore predictions for the SMBH bolometric luminosity function. Due to our conservative completeness limit there are only a relatively small number of objects with luminosities above this limit, resulting in noisier predictions than the SMBH mass function. Nevertheless, this analysis reveals a clear evolution from $z=10\to 5$ with the density of SMBHs with $L_{\rm bol}>10^{45}\ {\rm erg\ s^{-1}}$ increasing by around a factor $100$.
In Figure 6, we also compare our predictions to two recent observational studies. First we compare against the “free” and “polished” variants of the quasar bolometric luminosity function presented in Shen et al. (2020) at $z=5$ and $z=6$. These luminosity functions are derived by converting multiple observations of monochromatic quasar luminosity functions in the UV, optical, hard X-ray, and mid-IR into bolometric luminosity functions and finding the best fit. In the “free” variant the fitting of $\phi_{\star}$ is left free at each redshift. However, at high-redshift there is considerable observational uncertainty. In the “polished” variant $\phi_{\star}(z)$ is assumed to be linear with observations at $z=0.4-3.0$ used to define the relationship between $z$ and $\phi_{\star}$. While both variants provide excellent agreement with the FLARES predictions at $z=5-6$ it is worth noting that $L_{\rm bol}=10^{45-46}\ {\rm erg\ s^{-1}}$ is at the limit of the Shen et al. (2020) constraints.
[…]
One observational possibility for this discrepancy is potential contamination from host galaxy light or cosmic variance due to the relatively small volumes surveyed by JWST so far. Both issues will be overcome with increasing sample sizes and spectroscopic completeness. In terms of a theoretical explanation, beyond a fundamental issue with the model, there is some modelling flexibility in our predictions. First, here we use the instantaneous accretion rates, however, if we use accretion rates averaged over longer timescales we can obtain different results. This is explored in more detail in Appendix A. While the bright-end ($>10^{45.5}\ {\rm erg\ s^{-1}}$) of the LF remains immune to the choice of timescale (see Figure 13) we observe that the faint end can be enhanced by averaging over a longer timescale, bringing our predictions up to the lower luminosity constraints of Kokorev et al. (2024) and Greene et al. (2023). Second, here we assume a radiative efficiency of $\epsilon_{\rm r}=0.1$; increasing this to $\epsilon_{\rm r}=0.2$, as done by some other simulations, would shift the bolometric luminosity function by 0.3 dex to higher luminosities, again improving the agreement with Kokorev et al. (2024) and Greene et al. (2023).
4 Relation to Host Galaxy Physical Properties
We now explore the correlations between SMBH properties and their hosts.
4.1 $M_{\bullet}-M_{\star}$ scaling relationship
Figure 7: The relationship between stellar mass ($M_{\star}$) and SMBH mass ($M_{\bullet}$) predicted by FLARES for galaxies at $z=5$. In the top-panel only FLARES predictions are shown with individual objects are colour-coded by their Eddington ratio $\lambda$. The diagonal lines denote fixed values of $M_{\bullet}/M_{\star}$. The lower-panel instead shows $M_{\bullet}/M_{\star}$ across $105<z<$ 10, but restricts to objects with $L_{\rm bol}>10^{44}\ {\rm eg\ s^{-1}}$, and also includes observations at $z\sim 5$ from Maiolino et al. (2023b), Harikane et al. (2023) and Kocevski et al. (2023).
We begin by exploring the correlation between SMBH mass and the stellar content of their host galaxies. The top-panel of Figure 7 shows the relationship between $M_{\bullet}$ and $M_{\star}$ predicted for galaxies at $z=5$. For SMBHs with masses $>10^{7}\ {\rm M_{\odot}}$, the ratio of stellar to SMBH mass ($M_{\star}/M_{\bullet})$ is mostly in the range $100-2000$. There exist only a small number of galaxies where the SMBH has grown to exceed 1 per cent of the stellar mass. However, there are a significant number of galaxies in which the SMBH has yet to grow significantly beyond the seed mass, including relatively massive galaxies (i.e. those with $>10^{10}\ {\rm M_{\odot}}$). Consequently, in FLARES a tight relationship has yet to develop by $z=5$.
4.1.1 Comparison with Observational Constraints
The top-panel of Figure 7 shows a comparison of the FLARES predictions with the local empirical relations of Kormendy & Ho (2013b) and Reines & Volonteri (2015). If restricted to resolved SMBHs the FLARES predictions closely match the Kormendy & Ho (2013b) trend but lie somewhat above the Reines & Volonteri (2015) relation.
With the advent of JWST, measurements have become possible to higher redshift and lower luminosity. Several studies (e.g. Kocevski et al., 2023; Harikane et al., 2023; Maiolino et al., 2023b) have now measured stellar and SMBH masses of galaxies at $z>5$, albeit with small samples. These observations are included in the lower-panel of Figure 7 where we limit the FLARES predictions to $L_{\rm bol}>10^{44}\ {\rm eg\ s^{-1}}$ to better align with the completeness of the observations (and also present $M_{\bullet}/M_{\star}$ instead of $M_{\bullet}$ on the $y$-axis). This reveals an overall good correspondence with the majority of observational constraints intersecting with the FLARES predictions. It is worth noting that the observational constraints extend to $M_{\star}<10^{9}\ {\rm M_{\odot}}$, and where they do, the $M_{\bullet}/M_{\star}$ ratios are high (see Maiolino et al., 2023b). Such objects are not predicted by FLARES due to the mass resolution.
4.2 Relative Contribution of SMBHs to the Bolometric Luminosities of Galaxies
Figure 8: Ratio of SMBH bolometric luminosity to that of the stellar component of FLARES galaxies at $z\in[5,10]$ as a function of the total bolometric luminosity. Points are coloured coded by their stellar mass with outlined points those with $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. The dark solid line denotes the median of this relation with the dashed line denoting where number of objects in the bin are below five. The solid horizontal line gives the 1:1 relation, i.e. where $L_{\star}=L_{\bullet}$. Note, because of the FLARES simulation strategy each galaxy can have a unique weight and thus the median line will not be the median of un-weighted objects. The shaded region denotes where $L_{\bullet}<10^{45}\ {\rm erg\ s^{-1}}$.
Figure 9: The fraction of galaxies in which $L_{\bullet}>L_{\star}$ (solid thick lines) and $L_{\bullet}>0.1\times L_{\star}$ (dashed thin lines) as a function of the total bolometric luminosity for $z=5\to 10$.
We next explore the relative contribution of SMBH emission to the total bolometric luminosity of galaxies. In Figure 8 we present the ratio of the SMBH to stellar bolometric luminosity as a function of the total bolometric luminosity. The bolometric luminosities of objects with ratios $>1$ are then dominated by emission from a SMBH. In Figure 9 we instead show the fraction of galaxies in which the SMBH provides $>50\%$ or $\gtrsim 10\%$ of the bolometric luminosity.
While there is a large amount of scatter in this relation there is a clear trend of an increase in the $L_{\bullet}/L_{\star}$ ratio with total bolometric luminosity. Objects with $L_{\rm bol}>10^{45.5}\ {\rm erg/s}$ on average have SMBH bolometric luminosities surpassing starlight. Below $L_{\rm bol}\sim 10^{45}\ {\rm erg/s}$ the average fractional contribution appears to drop significantly. However, this may reflect the incompleteness in the black hole population in FLARES due to the resolution and modelling choices. From Figure 9 it is also clear that the relative contribution of SMBHs at fixed total bolometric luminosity drops with decreasing redshift. For galaxies with $L\sim 10^{45.5}\ {\rm erg\ s^{-1}}$ virtually all objects are SMBH dominated at $z=10$, dropping to $\approx 30\%$ at $z=5$.
This means that FLARES, unlike EAGLE, is exploring the regime in which emission from SMBHs is dominating the bolometric emission in galaxies. However, FLARES is only just doing this, and only simulates a handful of SMBHs with bolometric luminosities $>10^{46}\ \rm erg\ s^{-1}\approx 2\times 10^{12}\ {\rm L_{\odot}}$.
4.3 Total Bolometric Luminosity Function
Figure 10: Bolometric luminosity function for stars, SMBHs, and galaxies (combined stars and SMBHs) at $z\in[5,10]$ predicted by FLARES. Dashed (dotted) lines give the stellar (AGN) contribution to the total galaxy luminosity function given in solid lines.
Building on this analysis, in Figure 10 we show the bolometric function of stars, SMBHs, and the total at $105<z<$ 10. As anticipated from Figure 8 the bolometric luminosity function is dominated by stars at $L_{\rm bol}<10^{45}\ {\rm erg/s}$ but becomes dominated by SMBHs at $L_{\rm bol}>10^{45.5}\ {\rm erg/s}$, providing a clear boost to the total bolometric luminosity function.
4.4 The impact of AGN on their host galaxies
In EAGLE/FLARES accretion on to SMBHs releases energy, heating neighbouring gas particles. The FLARES simulation strategy makes it easy to experiment with changes to the model. To study the impact of AGN on galaxy formation, we re-simulate one of the high-density regions (region 03, $\delta_{14}\approx-0.31$) but with AGN feedback (see §2.2.6) turned off. In this variant SMBHs still grow, but they do not inject energy to the ISM.
In Figure 11 we show the difference between the mean specific star formation rate in bins of stellar mass. At the highest redshifts this is predictably noisy due to the small number of galaxies in this single simulation. At lower redshift however the number of galaxies has increased enough for us to be confident. This reveals that at low masses ($<10^{9}\ {\rm M_{\odot}}$) there is little or no impact on the average specific star formation rates of galaxies. However, there is tentative evidence for a suppression of star formation in the most massive ($>10^{10}\ {\rm M_{\odot}}$) galaxies due to the effect of AGN feedback.
Figure 11: The difference between the mean specific star formation rate with and without AGN feedback activated for galaxies $z=5-10$ for a single simulated FLARES region.
Lovell et al. (2023) studied the emergence of passive galaxies in the early Universe using FLARES. The EAGLE model produces number densities of passive galaxies in good agreement with observational constraints at $z<5$ (e.g Merlin et al., 2019; Carnall et al., 2023), which gives us confidence in looking at the passive populations at higher redshift. The main finding in Lovell et al. (2023) was that AGN feedback in particular was necessary to produce passive galaxies at $z\geqslant 5$, in agreement with the overall trends in specific star formation rate shown in Figure 11. Passive galaxies in FLARES are always those that have the largest SMBHs for their given stellar mass, however, while the growth of SMBHs was found to explain the suppression of star formation in passive galaxies, a large or accreting SMBH doesn’t necessarily result in the formation of a passive galaxy. Further investigating their SMBH accretion and star formation histories Lovell et al. (2023) found that the star formation activity in passive galaxies was anti-correlated with the SMBH accretion rate; passivity tended to follow a period of black hole accretion, and could persist for up to $\sim$400 Myr.
5 Conclusions
In this work, we have explored the physical and photometric properties of super-massive black holes (SMBHs) at high-redshift ($5\leq z\leq 10$) using the First Light And Reionisation Epoch Simulations (FLARES). FLARES is a suite of hydrodynamical zoom simulations probing a wide range of environments $\delta_{14}\approx-0.3\to 0.3$ making it ideally suited to studying rare objects. FLARES employs the well tested EAGLE physics model with the novel re-simulation strategy and post-processing to predict observational quantities. We predict the physical properties of SMBHs and their host galaxies and, through simulations without AGN feedback, we explore the effect of AGN on their host galaxies. Our conclusions are:
  • The resolution and modelling choices of FLARES limit us to the study of SMBHs with $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$. Compared to the $(100\ {\rm Mpc})^{3}$ EAGLE reference volume, FLARES simulates $\approx 8-20$ times more $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs at $z=5-9$ and $\approx 25$ times more $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ SMBHs at $z=6$, with samples of $M_{\bullet}>10^{8}\ {\rm M_{\odot}}$ SMBHs extending to $z=9$.
  • The number density of SMBHs predicted by FLARES drops by 10$\approx$ 10$\times$ 10 × from $z=5\to 7$. This trend may continue to higher-redshift but beyond $z=8$ the number of SMBHs simulated by FLARES is small. The density of SMBHs also drops by 10$\approx$ 10$\times$ 10 × for SMBHs with $M_{\bullet}=10^{7}\to 10^{8}\ {\rm M_{\odot}}$. FLARES predictions are compatible with recent observations of the $z=5$ broad line SMBH mass function; however there is significant uncertainty about the required completeness correction required to convert the broad line mass function to a total SMBH mass function. Where they overlap (in mass and redshift) FLARES is in good agreement with some other models, including Bluetides, Illustris, and TNG100, but lies above models including Astrid, Simba, and TNG300.
  • SMBHs are preferentially found in over-dense environments. The densest regions simulated by FLARES ($\delta_{14}=0.3$) have a density of $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs 20$\approx$ 20$\times$ 20 × higher than mean density regions. One implication is that observational studies will be strongly affected by cosmic variance.
  • At fixed mass, SMBHs in FLARES exhibit a range of accretion rates, with almost all having $10^{-6}\leq\lambda\leq 1$. The median $\lambda$ is $\approx 10^{-2}$, though decreases with $M_{\bullet}$.
  • The predicted SMBH bolometric luminosity evolves by $\sim 100$ from $z=10\to 5$. At $z=5-6$ it provides a close match to pre-JWST constraints. While the $z=5$ luminosity function is marginally consistent with JWST observations, at higher redshift there is increasing disagreement, particularly at the bright end. Some of this disagreement can be ameliorated by modelling choices, i.e. the choice of radiative efficiency and averaging timescale.
  • $M_{\bullet}>10^{7}\ {\rm M_{\odot}}$ SMBHs predominantly lie in galaxies with $M_{\star}>10^{9.5}\ {\rm M_{\odot}}$ and have masses $0.05-1.0\%$ of the stellar mass, comparable to local SMBHs. In FLARES no SMBH exceeds more than 2% of the stellar mass content in a galaxy. For galaxies hosting SMBHs with $L_{\rm bol}>10^{44}\ {\rm erg\ s^{-1}}$ the predicted $M_{\bullet}/M_{\star}$ are well matched to recent JWST observations.
  • The contribution of SMBHs to the bolometric luminosities of galaxies is found to rapidly increase as a function total bolometric luminosity. However, some of this increase may reflect incompleteness due to the resolution and modelling choices. For galaxies with $L_{\rm bol}>10^{45.5}\,{\rm erg\ s^{-1}}$ we find that accretion on to SMBHs, on average, dominates the total bolometric luminosity. This is reflected in the impact of SMBHs to the bolometric luminosity function (LF) of galaxies with SMBHs dominating at $L_{\rm bol}>10^{45.5}\,{\rm erg\ s^{-1}}$.
  • Using a pair of re-simulations of FLARES regions without AGN feedback enabled, we explore the impact of AGN on their host galaxies. By simply comparing the correlation between the stellar mass and specific star formation rate we find that AGN feedback has the effect of reducing the average star formation activity, but only in the most massive galaxies at the lowest redshifts explored by FLARES ($z=5-6$). In a companion work (Lovell et al., 2023) we explored the origin of passive galaxies predicted by FLARES finding that their passivity was driven by AGN feedback.
The advent of JWST has provided new observational tools to study galaxy formation and evolution in the distant Universe, particularly the role and contribution of SMBHs. While we have only begun scratching the surface of JWST’s contributions, there are already indications of tension between these observations and the FLARES predictions suggesting improvements to the physical model are required.
Looking to the future, JWST will continue building up observations - both larger imaging surveys, allowing us to photometrically identify AGN, and spectroscopic follow-up providing the means to unambiguously determine the contribute of AGN in composite objects. Furthermore, the Euclid spacecraft is now embarking on its mission to map almost the entire extragalactic sky in the optical and near-IR. These observations will allow us to identify bright/rare candidate AGN for subsequent follow-up by JWST.
While FLARES extends the range of masses and luminosities probed by hydrodynamical models with similar mass resolution it is still unable to reach the limit of observational studies. A second phase of FLARES is now underway which will both increase the simulated volume and resolution but also explore a wider set of models and model variations.
Changes from Version 1
In Version 1 of this manuscript we used the number densities presented in the original submission of Kokorev et al. (2024) in Figure 10. It was subsequently discovered that, due to an analysis error, these values were erroneously high by $\approx 0.2$ dex. In this version we use updated values from Kokorev et al. (2024). This change improves the agreement between our predictions and Kokorev et al. (2024), though a discrepancy still exists at $z>5$, particularly in the most luminous bins.
Author Contributions
We list here the roles and contributions of the authors according to the Contributor Roles Taxonomy (CRediT) [^2] https://credit.niso.org/. Jussi K. Kuusisto, Stephen M. Wilkins: Conceptualization, Data curation, Methodology, Investigation, Formal Analysis, Visualization, Writing - original draft. Christopher C. Lovell: Conceptualization, Data curation, Methodology, Writing - original draft. Dimitrios Irodotou, Shihong Liao, Sonja Soininen: Investigation, Writing - original draft. William Roper, Aswin P. Vijayan: Data curation, review & editing. Peter A. Thomas Conceptualization, Writing - review & editing. Sabrina C. Berger, Sophie L. Newman, Louise T. C. Seeyave, Shihong Liao: Writing - review & editing.
Acknowledgements
We thank the EAGLE  team for their efforts in developing the EAGLE  simulation code. This work used the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/K00042X/1, ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
JKK and LTCS are supported by an STFC studentship. SMW and WJR acknowledge support from the Sussex Astronomy Centre STFC Consolidated Grant (ST/X001040/1). CCL acknowledges support from a Dennis Sciama fellowship funded by the University of Portsmouth for the Institute of Cosmology and Gravitation. DI acknowledges support by the European Research Council via ERC Consolidator Grant KETJU (no. 818930) and the CSC – IT Center for Science, Finland. APV acknowledges support from the Carlsberg Foundation (grant no CF20-0534). The Cosmic Dawn Center (DAWN) is funded by the Danish National Research Foundation under grant No. 140. SL acknowledges the supports by the National Natural Science Foundation of China (NSFC) grant (No. 11988101) and the K. C. Wong Education Foundation.
We also wish to acknowledge the following open source software packages used in the analysis: Numpy (Harris et al., 2020), Scipy (Virtanen et al., 2020), Astropy (Astropy Collaboration et al., 2013, 2018, 2022), Cmasher (van der Velden, 2020), and Matplotlib (Hunter, 2007).
Data Availability Statement
The data associated with the paper will be made publicly available at https://flaresimulations.github.io on the acceptance of the manuscript.
Appendix A Instantaneous vs. time averaged accretion rates
In this work we have utilised the instantaneous accretion rates. However, since, in the FLARES/EAGLE model these rates can vary dramatically between time-steps it is interesting to explore how averaging on various timescales affects some of our results.
We being in Figure 12 by showing the ratio of the instantaneous accretion rate to the accretion rate averaged over the preceding 50 Myr for SMBHs at $z=5$. This reveals that the majority of galaxies have average accretion rates larger than their instantaneous rate. Despite this, the total amount of accretion, across all black holes, is actually larger, albeit only slightly ($<10\%$), for instantaneous accretion rates.
Figure 12: The ratio of the instantaneous accretion rate to the accretion rate averaged over the preceding 50 Myr for SMBHs at $z=5$. The solid line denoted the median in $M_{\bullet}$ bins. Individual BHs are colour coded by their Eddington ratio ($\lambda$).
This effect has important implications for the accretion rate distribution function, or bolometric luminosity function. Figure 13 shows the bolometric luminosity function calculated assuming the different accretion timescales $\in\{10,20,50,100,200\}\ {\rm Myr}$. While the density of very-luminous ($>10^{46}\ {\rm erg\ s^{-1}}$) SMBHs remains relatively unchanged, the number of lower luminosity, but still bright ($>10^{44.5-45.5}\ {\rm erg\ s^{-1}}$) SMBHs increases significantly. This is due to many BHs with very low instantaneous accretion rates having significantly higher averaged rates as demonstrated by Figure 12.
Figure 13: The bolometric luminosity function of SMBHs using both the instantaneous accretion rates (thick black line) and varying averaging timescales $\in\{10,20,50,100,200\}\ {\rm Myr}$.